OpenStudy (anonymous):

log (x+3) + log x = 2

OpenStudy (mathstudent55):

Use the rule of logs: \( \log a + \log b = \log (ab) \)

OpenStudy (solomonzelman):

a= x+3 b= x use the rule that student wrote.

OpenStudy (mathstudent55):

Then use the definition of log: \( \log_{10} x = y\) means \( 10^y = x\)